Assume that a 2 kg of a certain gas with R=0.4305 kJ/kg-K undergoes a process that results in these changes: ΔU=1458.01 kJ, ΔH=2189.13 kJ. If the process had been internally reversible with p=C, what is the work done?
1
Expert's answer
2021-07-12T03:37:08-0400

"C_p-C_v = R" ...(1)


"k = \frac{C_p}{C_v}" ...(2)


From 1 and 2

"R = 1.49C_v-C_v = 0.49C_v"


"C_v = \frac{0.4305}{0.49}=0.8785" "\frac {kJ} {kg K}"


"Q = mC_v\triangle U"


"Q =2 \times 0.8785 \times 1458.01 = 2561.72 kJ"


work done "W = Q - \triangle U =2561.72 - 1458.01 =1103.71 kJ"